# AGALE: A Graph-Aware Continual Learning Evaluation Framework

### v1

@@ Abstract
In recent years, continual learning (CL) techniques have made significant progress in learning from streaming data while preserving knowledge across sequential tasks, particularly in the realm of euclidean data. To foster fair evaluation and recognize challenges in CL settings, several evaluation frameworks have been proposed, focusing mainly on the single- and multi-label classification task on euclidean data. However, these evaluation frameworks are not trivially applicable when the input data is graph-structured, as they do not consider the topological structure inherent in graphs. Existing continual graph learning (CGL) evaluation frameworks have predominantly focussed on single-label scenarios in the node classification (NC) task. This focus has overlooked the complexities of multi-label scenarios, where nodes may exhibit affiliations with multiple labels, simultaneously participating in multiple tasks. We develop a graph-aware evaluation (AGALE) framework that accommodates both single-labeled and multi-labeled nodes, addressing the limitations of previous evaluation frameworks. In particular, we define new incremental settings and devise data partitioning algorithms tailored to CGL datasets. We perform extensive experiments comparing methods from the domains of continual learning, continual graph learning, and dynamic graph learning (DGL). We theoretically analyze AGALE and provide new insights about the role of homophily in the performance of compared methods. We release our framework at https://github.com/Tianqi-py/AGALE.
@@ 1.1 Limitations of current continual learning evaluation frameworks
The multi-label continual learning evaluation frameworks for Euclidean data (Kim et al., 2020a) suggest the use of hierarchical clustering techniques, which involves grouping classes into a single task based on their co-occurrence frequency and subsequently eliminating instances with label sets that do not align with any class groups. Applying such a technique to graph-structured data not only risks excluding nodes with a higher number of labels but also disrupts the associated topological structures.
@@
Last but not least, previous CGL benchmarks typically predetermined the model’s output units, assuming an approximate count of classes in each graph during model initialization. However, this assumption is unrealistic because the eventual class size is often unknown, leading to potential inefficiencies in storage or capacity overflow.
@@ 1.2 Our Contributions
To tackle the above-mentioned gaps, we develop a generalized evaluation framework for continual graph learning, which is applicable both for multi-class and multi-label node classification tasks and can be easily adapted for multi-label graph and edge classification tasks. To summarize, our contributions are as follows.
@@ 3.2 Data Partitioning Algorithms
  • Task sequence and subgraph sequence generation. We employ Algorithm 1 to segment the provided graph from the dataset into coherent subgraph sequences.
    Instead of using a predefined class order (as discussed in Section 1.1) we generate $P$ random orders of the classes to simulate the random emergence of new elements in the real world. For multi-label datasets, the random orders also allow dynamic correlation between the classes in the incremental settings. Specifically, given a dataset with $C$ classes, we group $K$ random classes as one task for one time step. At any time point $t$, let $\mathcal{C}_{t}$ denote the set of classes grouped for the task at time $t$. We construct a subgraph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ such that $\mathcal{V}_{t}$ is the set of nodes with one or more labels in $\mathcal{C}_{t}$. The edge set $\mathcal{E}_{t}$ consists of the induced edges on $\mathcal{V}_{t}$. Note that the number of classes chosen to create one task is adaptable. In order to maximize the length of the task sequence for each given graph dataset and subsequently catastrophic forgetting, we choose $K=2$ in this work. If the dataset has an uneven number of classes in total, the remaining last class will be grouped with the second last class group.
@@ A.3 Time And Space Complexity Analysis
For GCN, at each layer, the operation includes feature transformation, neighbourhood aggregation, and activation. The feature transformation over two layers leads to the multiplication of matrices of sizes (i) $n\times d$ and $d\times d$, and (ii) $d\times d$ and $d\times d$ which leads to a total time complexity of $O(nd^{2})$.
And the neighborhood aggregation requires a multiplication between matrices of size $n\times n$ and $n\times d$, yielding $O(n^{2}d)$. In practice, we compute this using a sparse operator, such as the PyTorch scatter function for each entry $(i,j)$ in the adjacency matrix of the edge $e\in\mathcal{E}$, which yields a total cost of $O(md)$. Finally, the activation is an element-wise function with the time complexity of $O(n)$. Overall, the time complexity of a $L$ layer GCN is $O(nd^{2}L+mdL+nL)$.
For computing space requirements of GCN, we include (i) the space required for the input adjacency matrix of size $n\times n$, (ii) the feature matrix of size $n\times d$, and (iii) the model itself with $d^{2}+d$ parameters for weight and bias in each layer. In total, the space complexity of GCN is $O(n^{2}+nd+L(d^{2}+d))$.
@@
SimpleGCN trains a GCN for each time step $t\in\mathcal{T}$. And because it does not apply any continual learning techniques to remember from previous time steps, the time is the same with GCN, i.e., $O(|\mathcal{T}|T_{GCN})$ and the space complexity is $S_{GCN}$.
@@
LwF uses GCN as the backbone model, the GCN is trained at each time step $t\in\mathcal{T}$ for the new task, which gives the time complexity of $O(\mathcal{|T|}T_{GCN})$. To do the knowledge distillation, the previous model also calls GCN forward passes with time complexity of $O(\mathcal{|T|}T_{GCN})$. Overall, the time complexity is $O(2\mathcal{|T|}T_{GCN})$. The space consumption consists of loading the current GCN model and the previous GCN model for the knowledge distillation, with space complexity of $O(2S_{GCN})$.
@@
EWC calls forward passes of GCN at each time step, and for each parameter at each time step $t\in\mathcal{T}$, the values on the diagonal of the fisher matrix are approximated using the value of each model parameter itself and its gradient. This is an element-wise calculation, which gives the time complexity of $O(\mathcal{|T|}\times M)$, $M$ indicates the number of parameters in the GCN, which is $L(d^{2}+d)$. Overall, the time complexity of EWC is the time complexity of GCN plus the calculation of the fisher matrix, which is $O(\mathcal{|T|}\times(T_{GCN}+M))$. The space requirement of EWC consists of the space requirement of GCN, and the matrix stores the gradients of the parameters at each time step of size $O(\mathcal{|T|}\times M)$. In total, the space complexity of EWC is $O(S_{GCN}+\mathcal{|T|}M)$.
@@
Similarly, MAS using GCN as backbone model, at each time step $t\in\mathcal{T}$, there are forward passes of GCN and the calculation of fisher matrix for parameters, which gives the overall time complexity of $O(\mathcal{|T|}(T_{GCN}+M))$. The space requirement of MAS consists of the space complexity of GCN and one matrix for the gradient of the parameters of size $O(M)$, which yields in total $O(S_{GCN}+M)$.
@@
EvolveGCN is a method from the category of Dynamic Graph Neural networks, which trains a new model at each time step with time complexity same as GCN, i.e., $T_{GCN}$ and updates the model parameter using a recurrent neural network using the corresponding parameter from previous time step as input, which has the time complexity of $O(|\mathcal{T}|nd)$. Overall, EvolveGCN is more expensive than the other Continual Learning methods with time complexity of $O(T_{GCN}+|\mathcal{T}|nd)$. The space requirement of EvolveGCN consists of the space requirement of the GCN plus the recurrent unit for the reset, update, and new gate with $3(d^{2}+d)$. In our implementation, we use one recurrent layer for each layer in GCN. Thus the overall space complexity yields $O(2S_{GCN}+3M)$.
@@
ERGNN is a replay-based method. At each time step $t\in\mathcal{T}$, it retrains the GCN on the current graph and the buffer-nodes-formed graph, and the sampling process goes through the nodes in the new graph, which gives the time complexity of $O(2|\mathcal{T}|T_{GCN}+n)$. The space requirement of ERGNN consists of the buffer size of $\mathcal{|B|}$ and the space complexity of GCN. In total, it yields the space complexity of $O(S_{GCN}+\mathcal{|B|})$.
@@
Table 8: The simplified time complexity analysis, especially the constant term and the lower order terms, are dropped.
@@
SimpleGCN $O(|\mathcal{T}|T_{GCN})$ $S_{GCN}$
LwF $O(2\mathcal{|T|}T_{GCN})$ $O(2S_{GCN})$
EWC $O(\mathcal{|T|}\times(T_{GCN}+M))$ $O(S_{GCN}+\mathcal{|T|}M)$
MAS $O(\mathcal{|T|}(T_{GCN}+M))$ $O(S_{GCN}+M)$
EvolveGCN $O(T_{GCN}+|\mathcal{T}|nd)$ $O(2S_{GCN}+3M)$
ERGNN $O(2|\mathcal{T}|T_{GCN}+n)$ $O(S_{GCN}+\mathcal{|B|})$

### v2

@@ Abstract
In recent years, continual learning (CL) techniques have made significant progress in learning from streaming data while preserving knowledge across sequential tasks, particularly in the realm of euclidean data. To foster fair evaluation and recognize challenges in CL settings, several evaluation frameworks have been proposed, focusing mainly on the single- and multi-label classification task on euclidean data. However, these evaluation frameworks are not trivially applicable when the input data is graph-structured, as they do not consider the topological structure inherent in graphs. Existing continual graph learning (CGL) evaluation frameworks have predominantly focused on single-label scenarios in the node classification (NC) task. This focus has overlooked the complexities of multi-label scenarios, where nodes may exhibit affiliations with multiple labels, simultaneously participating in multiple tasks. We develop a graph-aware evaluation (AGALE) framework that accommodates both single-labeled and multi-labeled nodes, addressing the limitations of previous evaluation frameworks. In particular, we define new incremental settings and devise data partitioning algorithms tailored to CGL datasets. We perform extensive experiments comparing methods from the domains of continual learning, continual graph learning, and dynamic graph learning (DGL). We theoretically analyze AGALE and provide new insights about the role of homophily in the performance of compared methods. We release our framework at https://github.com/Tianqi-py/AGALE.
@@ 1.1 Limitations of current continual learning evaluation frameworks
The multi-label continual learning evaluation framework for Euclidean data (Kim et al., 2020a) suggest the use of hierarchical clustering techniques, which involves grouping classes into a single task based on their co-occurrence frequency and subsequently eliminating instances with label sets that do not align with any class groups. Applying such a technique to graph-structured data not only risks excluding nodes with a higher number of labels but also disrupts the associated topological structures.
@@
Last but not least, previous CGL benchmarks typically predetermined the size of model’s output units, assuming an approximate count of classes in each graph during model initialization. However, this assumption is unrealistic because the eventual class size is often unknown, leading to potential inefficiencies in storage or capacity overflow.
@@ 1.2 Our Contributions
To tackle the above-mentioned gaps, we develop a generalized evaluation framework for continual graph learning, which is applicable both for multi-class and multi-label node classification tasks and can be easily adapted for multi-label graph and edge classification tasks. Our key contributions are as follows.
@@ 3.2 Data Partitioning Algorithms
  • Task sequence and subgraph sequence generation. We employ Algorithm 1 to segment the provided graph from the dataset into coherent subgraph sequences. We first remove the classes with size smaller than a threshold $\delta$. Instead of using a predefined class order (as discussed in Section 1.1) we generate $n$ random orders of the remaining classes to simulate the random emergence of new classes in the real world. Specifically, given a dataset with $C$ classes, we group $K$ random classes as one task for one time step. At any time point $t$, let $\mathcal{C}_{t}$ denote the set of classes grouped for the task at time $t$. We construct a subgraph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ such that $\mathcal{V}_{t}$ is the set of nodes with one or more labels in $\mathcal{C}_{t}$. The edge set $\mathcal{E}_{t}$ consists of the induced edges on $\mathcal{V}_{t}$. Note that the number of classes chosen to create one task is adaptable. In order to maximize the length of the task sequence for each given graph dataset and subsequently catastrophic forgetting, we choose $K=2$ in this work. If the dataset has an uneven number of classes in total, the remaining last class will be grouped with the second last class group.
@@ A.3 Time And Space Complexity Analysis
For GCN, at each layer, the operation includes feature transformation, neighbourhood aggregation, and activation. The feature transformation over two layers leads to the multiplication of matrices of sizes (i) $n\times d$ and $d\times d$, and (ii) $d\times d$ and $d\times d$ which leads to a total time complexity of $\mathcal{O}(nd^{2})$.
And the neighborhood aggregation requires a multiplication between matrices of size $n\times n$ and $n\times d$, yielding $\mathcal{O}(n^{2}d)$. In practice, we compute this using a sparse operator, such as the PyTorch scatter function for each entry $(i,j)$ in the adjacency matrix of the edge $e\in\mathcal{E}$, which yields a total cost of $\mathcal{O}(md)$. Finally, the activation is an element-wise function with the time complexity of $\mathcal{O}(n)$. Overall, the time complexity of a $L$ layer GCN is $\mathcal{O}(nd^{2}L+mdL+nL)$.
For computing space requirements of GCN, we include (i) the space required for the input adjacency matrix of size $n\times n$, (ii) the feature matrix of size $n\times d$, and (iii) the model itself with $d^{2}+d$ parameters for weight and bias in each layer. In total, the space complexity of GCN is $\mathcal{O}(n^{2}+nd+L(d^{2}+d))$.
@@
SimpleGCN trains a GCN for each time step $t\in\mathcal{T}$. And because it does not apply any continual learning techniques to remember from previous time steps, the time is the same with GCN, i.e., $\mathcal{O}(|\mathcal{T}|T_{GCN})$ and the space complexity is $S_{GCN}$.
@@
LwF uses GCN as the backbone model, the GCN is trained at each time step $t\in\mathcal{T}$ for the new task, which gives the time complexity of $\mathcal{O}(\mathcal{|T|}T_{GCN})$. To do the knowledge distillation, the previous model also calls GCN forward passes with time complexity of $\mathcal{O}(\mathcal{|T|}T_{GCN})$. Overall, the time complexity is $\mathcal{O}(2\mathcal{|T|}T_{GCN})$. The space consumption consists of loading the current GCN model and the previous GCN model for the knowledge distillation, with space complexity of $\mathcal{O}(2S_{GCN})$.
@@
EWC calls forward passes of GCN at each time step, and for each parameter at each time step $t\in\mathcal{T}$, the values on the diagonal of the fisher matrix are approximated using the value of each model parameter itself and its gradient. This is an element-wise calculation, which gives the time complexity of $\mathcal{O}(\mathcal{|T|}\times M)$, $M$ indicates the number of parameters in the GCN, which is $L(d^{2}+d)$. Overall, the time complexity of EWC is the time complexity of GCN plus the calculation of the fisher matrix, which is $\mathcal{O}(\mathcal{|T|}\times(T_{GCN}+M))$. The space requirement of EWC consists of the space requirement of GCN, and the matrix stores the gradients of the parameters at each time step of size $\mathcal{O}(\mathcal{|T|}\times M)$. In total, the space complexity of EWC is $\mathcal{O}(S_{GCN}+\mathcal{|T|}M)$.
@@
Similarly, MAS using GCN as backbone model, at each time step $t\in\mathcal{T}$, there are forward passes of GCN and the calculation of fisher matrix for parameters, which gives the overall time complexity of $\mathcal{O}(\mathcal{|T|}(T_{GCN}+M))$. The space requirement of MAS consists of the space complexity of GCN and one matrix for the gradient of the parameters of size $\mathcal{O}(M)$, which yields in total $\mathcal{O}(S_{GCN}+M)$.
@@
EvolveGCN is a method from the category of Dynamic Graph Neural networks, which trains a new model at each time step with time complexity same as GCN, i.e., $T_{GCN}$ and updates the model parameter using a recurrent neural network using the corresponding parameter from previous time step as input, which has the time complexity of $\mathcal{O}(|\mathcal{T}|nd)$. Overall, EvolveGCN is more expensive than the other Continual Learning methods with time complexity of $\mathcal{O}(T_{GCN}+|\mathcal{T}|nd)$. The space requirement of EvolveGCN consists of the space requirement of the GCN plus the recurrent unit for the reset, update, and new gate with $3(d^{2}+d)$. In our implementation, we use one recurrent layer for each layer in GCN. Thus the overall space complexity yields $\mathcal{O}(2S_{GCN}+3M)$.
@@
ERGNN is a replay-based method. At each time step $t\in\mathcal{T}$, it retrains the GCN on the current graph and the buffer-nodes-formed graph, and the sampling process goes through the nodes in the new graph, which gives the time complexity of $\mathcal{O}(2|\mathcal{T}|T_{GCN}+n)$. The space requirement of ERGNN consists of the buffer size of $\mathcal{|B|}$ and the space complexity of GCN. In total, it yields the space complexity of $\mathcal{O}(S_{GCN}+\mathcal{|B|})$.
@@
Table 8: The simplified time complexity analysis. $T_{GCN}$ and $S_{GCN}$ corresponds to the time and space requirement of the base GCN model.
@@
SimpleGCN $\mathcal{O}(|\mathcal{T}|T_{GCN})$ $S_{GCN}$
LwF $\mathcal{O}(2\mathcal{|T|}T_{GCN})$ $\mathcal{O}(2S_{GCN})$
EWC $\mathcal{O}(\mathcal{|T|}\times(T_{GCN}+M))$ $\mathcal{O}(S_{GCN}+\mathcal{|T|}M)$
MAS $\mathcal{O}(\mathcal{|T|}(T_{GCN}+M))$ $\mathcal{O}(S_{GCN}+M)$
EvolveGCN $\mathcal{O}(T_{GCN}+|\mathcal{T}|nd)$ $\mathcal{O}(2S_{GCN}+3M)$
ERGNN $\mathcal{O}(2|\mathcal{T}|T_{GCN}+n)$ $\mathcal{O}(S_{GCN}+\mathcal{|B|})$